OpenStudy (anonymous):

Solve the problem by using substitution. Be sure to show all your work. x+y=5 y=x-1

OpenStudy (popitree):

x+(x-1)=5 >2x=6 >x=3 >y=3-1=2

OpenStudy (anonymous):